# Slant sums of quiver gauge theories

### v1

@@ Abstract
We define a notion of slant sum of quiver gauge theories, a type of surgery on the underlying quiver. Under some mild assumptions, we relate torus fixed points on the corresponding Higgs branches, which are Nakajima quiver varieties. Then we prove a formula relating the quasimap vertex functions before and after a slant sum, which is a type of “branching rule” for vertex functions.
Our construction is motivated by a conjecture, which we make here, for the factorization of the vertex functions of zero-dimensional quiver varieties. The branching rule allows this conjecture to be approached inductively. In special cases, it also provides a formula for the $\hbar=q$ specialization of vertex functions for quiver varieties not necessarily of Dynkin type as a sum over reverse plane partitions. When passed through the quantum Hikita conjecture, such expressions provide conjectural formulas for graded traces of Verma modules on the 3d mirror dual side.
We also consider the Coulomb side. We make some conjectures reflecting what can be seen on the Higgs side and prove them in ADE type. We study slant sums of Coulomb branches and their quantizations. We prove that for one-dimensional framing, slant sum operation on the Coulomb branch side corresponds to taking products.
@@ 1.1 Slant sums of heaps
The slant sum of two posets was defined by Proctor in [38] to break certain posets into simpler, slant irreducible pieces. Some particular posets to which this applies are those that appear as heaps $H(w)$ of fully commutative elements $w$ of a Weyl group $W$, see [39]. Such posets are defined in a straightforward combinatorial way from a choice of reduced word for $w$.
It was noticed in [18] that the heap of a minuscule element gives rise to a module over the preprojective algebra of the corresponding Dynkin quiver. Equivalently, it gives rise to a point in a Nakajima quiver variety. When $w$ is dominant minuscule, the corresponding quiver variety is a single point, and the preprojective algebra module studied in [18] is a choice of representative of this point.
Suppose $w$, $w^{(1)}$, and $w^{(2)}$ are dominant minuscule elements of some Weyl groups and let $\mathcal{M}$, $\mathcal{M}^{(1)}$, and $\mathcal{M}^{(2)}$ be the respective zero-dimensional Nakajima quiver varieties. The claim that the heap of $w$ decomposes as a slant sum of the other two heaps, written $H(w)=H(w^{(1)})\#H(w^{(2)})$, is equivalent to saying that a representative of the unique point in $\mathcal{M}$ can be built from representatives of the unique points in $\mathcal{M}^{(1)}$ and $\mathcal{M}^{(2)}$. It is thus reasonable to call $\mathcal{M}$ the slant sum of the two other quiver varieties.
The aim of this paper is to revisit slant sums from the perspective of quiver varieties, or more precisely, quiver gauge theories. If the numerical data used to construct $\mathcal{M}^{(1)}$ and $\mathcal{M}^{(2)}$, which now need not be zero-dimensional, is compatible, we define a new quiver variety $\mathcal{M}:=\mathcal{M}^{(1)}\#\mathcal{M}^{(2)}$ which we call the slant sum of the two constituents. Unlike in [18], our procedure quickly leaves the world of Dynkin quivers, which we view as a feature.
@@ 1.2 Slant sums of quiver gauge theories
Let $Q^{(r)}$, $\mathsf{v}^{(r)}$, and $\mathsf{w}^{(r)}$ for $r\in\{1,2\}$ be two such collections. Suppose there exist vertices $\star_{1}\in Q^{(1)}_{0}$ and $\star_{2}\in Q^{(2)}_{0}$ such that $\mathsf{v}^{(1)}_{\star_{1}}=\mathsf{w}^{(2)}_{\star_{2}}$. We say that the two quiver gauge theories are slant summable, and we define their slant sum to be the quiver gauge theory defined by identifying the gauge vertex at $\star_{1}$ with the framing vertex at $\star_{2}$. We will use the symbol $_{\star_{1}}\#_{\star_{2}}$, or just #\#, throughout to denote the effect of this operation on various objects. For example, the quiver for the slant sum of the two quiver gauge theories is $Q=Q^{(1)}{}_{\star_{1}}\#_{\star_{2}}Q^{(2)}$. It is obtained by adding a single arrow to the disjoint union of the two constituent quivers.
@@
In the setting of the previous proposition, we will also denote $p^{(1)}\#p^{(2)}:=\Psi_{p^{(1)}}(p^{(2)})$. When $\mathcal{M}$, $\mathcal{M}^{(1)}$, and $\mathcal{M}^{(2)}$ are 0-dimensional quiver varieties corresponding to dominant minuscule Weyl group elements and $\mathsf{v}^{(1)}_{\star_{1}}=\mathsf{w}^{(2)}_{\star_{2}}=1$, Theorem 1.1 recovers the classical notion of slant sums of heaps studied in [38], see also [18].
@@ 1.3 Slant sums and vertex functions
Our constructions also relate to certain quasimap counts. In the $K$-theoretic enumerative geometry of quiver varieties, the descendant vertex function is a key object [37]. It appears in Okounkov’s enumerative 3d mirror symmetry [1, 36], see also [5, 12, 13]. It is defined as the generating function of equivariant counts of quasimaps from $\mathbb{P}^{1}$ to $\mathcal{M}$ with a nonsingular condition at $\infty\in\mathbb{P}^{1}$. We denote it by
@@
and it is natural to wonder if there is any relationship between $V^{(\tau)}_{p}$, $V^{(\tau_{1})}_{p^{(1)}}$, and $V^{(\tau_{2})}_{p^{(2)}}$ when $\tau=\iota^{*}(\tau_{1}\otimes\tau_{2})$. This is the content of our main theorem.
@@
Assume the setting of Theorem $1.$ Let $\tau=\iota^{*}(\tau_{1}\otimes\tau_{2})\in K_{\mathsf{T}\times G_{\mathsf{v}}}(\mathrm{pt})$. Then
@@
where the sum runs over $\mathsf{T}^{(1)}\times\mathbb{C}^{\times}_{q}$-fixed components $F$ of $\mathsf{QM}_{p^{(1)}}$
@@
We explore several corollaries in Section 3.4 in which Theorem 3.2 can be made more transparent. In the special case when $\mathcal{M}^{(2)}$ is zero-dimensional, $V^{(1)}_{\mathcal{M}^{(2)}}$ will not depend on the so-called framing parameters. It follows that the twisted vertex is equal to the ordinary vertex up to a monomial in $z_{2}$, and we obtain the factorization
$V^{(\tau_{1}\otimes 1)}_{p}(z_{1},z_{2})=V^{(\tau_{1})}_{p^{(1)}}(z_{1}^{\prime})V^{(1)}_{p^{(2)}}(z_{2})$
where $z_{1}^{\prime}$ stands for a shift of $z_{1}$ by some powers of $z_{2}$, see Corollary 3.5.
1.4 Connection with quantum Hikita conjecture
As another special case, we can take the $q=\hbar$ specialization of Theorem 1.2, where $\hbar$ is the weight of the symplectic form of $\mathcal{M}$ and $q$ is the equivariant parameter for the action of $\mathbb{C}^{\times}$ on the source $\mathbb{P}^{1}$ of quasimaps. For certain descendants, this again forces each twisted vertex in Theorem 1.2 to be equal to the untwisted vertex up to a monomial in $z_{2}$. If we are lucky, this monomial can be absorbed into $z_{1}$, leading to another factorization
$V^{(\tau_{1}\otimes\tau_{2})}_{p}(z_{1},z_{2})\big|_{q=\hbar}=V^{(\tau_{1})}_{p^{(1)}}(z_{1}^{\prime})\big|_{q=\hbar}V^{(\tau_{2})}_{p^{(2)}}(z_{2})\big|_{q=\hbar},$ (2)
see Corollary 3.7 for the precise statement.
This is related to the quantum Hikita conjecture proposed in [22], which predicts an isomorphism between the $q=\hbar$ specialized quantum $D$-module for the Higgs branch $\mathcal{M}$ and the $D$-module of graded traces for the Coulomb branch.
A proof of this conjecture was given for ADE quiver varieties with minuscule framings in [15]. The technical heart of the proof is to identify the $q=\hbar$ specialized descendant vertex functions with the graded traces of elements of the quantized Coulomb branch–in this case a truncated shifted Yangian–over a Verma module. Thus our results here provide many cases of the Higgs side of the computation needed to prove the quantum Hikita for more general quiver varieties. In many examples, see Section 4.1, vertex functions can be computed in terms of reverse plane partitions over some poset. Theorem 1.2 shows that this is true for $V^{(\tau)}_{p}$ if it is true for $V^{(\tau_{1})}_{p^{(1)}}$ and $V^{(\tau_{2})}_{p^{(2)}}$.
Furthermore, passing (2), or more generally Theorem 1.2, through 3d mirror symmetry gives a conjectural “branching rule” for graded traces. To the best of the authors’ knowledge, this formula, and its expression in terms of reverse plane partitions in special cases, is a new expectation.
1.5 Conjecture for 0-dimensional vertex functions
Theorem 1.2 and our study of slant sums was motivated by 3d mirror symmetry. Let $\mathsf{A}\subset\mathsf{T}$ be the subtorus preserving the symplectic form and let $p\in\mathcal{M}^{\mathsf{A}}$. As proposed in Conjecture 1 in [16], the limit in the equivariant parameters $\mathsf{A}$ of the vertex $V_{p}$ should factorize to a product of $q$-binomials, one for each repelling weight of $T_{p^{!}}\mathcal{M}^{!}$, the tangent space of the mirror dual variety $\mathcal{M}^{!}$ at the dual fixed point $p^{!}$ [^2] It is also expected, and assumed here, that there is a bijection between fixed points on $\mathcal{M}$ and $\mathcal{M}^{!}$ when all fixed points are isolated.. This is a degeneration of the full 3d mirror symmetry of vertex functions proven in type $A$ in [5]. Conveniently, it can be checked even when the definition of vertex function of $\mathcal{M}^{!}$ is unknown, which is currently the case except in type $A$.
In particular, if $\mathcal{M}$ is a single point, the vertex function $V_{\mathcal{M}}$ is already independent of the equivariant parameters of $\mathsf{A}$ and is thus expected to factorize to a product of $q$-binomials. Here we make two conjectures refining the conjecture of [16], providing an explicit formula for the factorization and, equivalently, for the tangent space of $\mathcal{M}^{!}$.
@@ 1.5 Conjecture for 0-dimensional vertex functions
To state it, we fix a quiver $Q$ without loops and a zero-dimensional quiver variety $\mathcal{M}:=\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})$ where the stabilty parameter is $\theta=(1,1,\ldots,1)$. As usual, we define weights
$\lambda:=\sum_{i\in Q_{0}}\mathsf{w}_{i}\varpi_{i},\quad\mu:=\lambda-\sum_{i\in Q_{0}}\mathsf{v}_{i}\alpha_{i}$ (3)
@@
of the Kac-Moody Lie algebra associated to the quiver $Q$. Here, $\varpi_{i}$ and $\alpha_{i}$ denote the fundamental weights and simple roots, respectively. Let $\Phi^{\pm,\text{re}}$ be the set of real positive or negative roots. Let $\Phi^{-,\text{re}}_{\mu}=\{\alpha\in\Phi^{-,\text{re}}\,\mid\,(\alpha,\mu)>0\}$ and $\Phi^{+,\text{re}}_{\mu}=\{\alpha\in\Phi^{+,\text{re}}\,\mid\,(\alpha,\mu)<0\}$. One can show, see Lemma 6.4, that $\dim\mathcal{M}=0\iff\mu\in W\lambda$.
@@
The vertex function depends on a choice of polarization, and we choose the canonical polarization [^3] The is determined by the choice of orientation of the edges in the underlying graph of the quiver. Equivalently, we choose the polarization induced by the zero section of the cotangent bundle appearing in the symplectic reduction defining $\mathcal{M}$.. Then we define
$\mathsf{V}_{\mathcal{M}}(z)=V_{\mathcal{M}}(z)_{z=z(-\hbar^{-1/2})^{\det T^{1/2}}}$ (4)
i.e. the shift is
@@
$z_{i}=z_{i}(-\hbar^{-1/2})^{a_{i}},\quad a_{i}=\sum_{\begin{subarray}{c}e\\ h(e)=i\end{subarray}}\mathsf{v}_{t(e)}-\sum_{\begin{subarray}{c}e\\ t(e)=i\end{subarray}}\mathsf{v}_{h(e)}+w_{i}$
@@
Let $A_{Q}=(a_{i,j})_{i,j\in Q_{0}}$ be the adjacency matrix of the quiver, i.e., $a_{i,j}$ is the number of arrows from vertex $i$ to vertex $j$. We identify the Kähler parameters with formal exponentials of simple roots via $e^{\alpha_{i}}=(q/\hbar)^{(\mathsf{v}-A_{Q}\cdot\mathsf{v})_{i}}z_{i}$. More explicitly,
$e^{\alpha_{i}}=z_{i}(q/\hbar)^{b_{i}},\quad b_{i}=\mathsf{v}_{i}-\sum_{\begin{subarray}{c}e\\ t(e)=i\end{subarray}}\mathsf{v}_{h(e)}$
Let $\Phi(x):=\prod_{i=0}^{\infty}(1-xq^{i})$.
@@
$\mathsf{V}_{\mathcal{M}}(z)=\prod_{\begin{subarray}{c}\alpha\in\Phi^{+,\text{re}}_{\mu}\\ \end{subarray}}\prod_{i=1}^{-(\alpha,\mu)}\frac{\Phi\left(\hbar\left(\frac{\hbar}{q}\right)^{i-1}e^{\alpha}\right)}{\Phi\left(\left(\frac{\hbar}{q}\right)^{i-1}e^{\alpha}\right)}$
This conjecture has been proven for type $A$ quiver varieties with one framing by Smirnov and the first author in [17]. It was also proven by Jang and the first author for type $D$ quiver varieties with a single minuscule framing in [14]. Finally, for cotangent bundles of Grassmannians, it follows from [13]. One of our motivations in studying slant sums of quiver varieties was to develop tools to prove this conjecture more generally.
Equivalently, under the identification $e^{\alpha_{i}}=z_{i}(q/\hbar)^{-b_{i}}$, we have
@@
In this formula, we now view $\mu$ as a coweight using the same formulas as (3) but with fundamental coweights and simple coroots replacing fundamental weights and simple roots. We now write the pairing as $\langle\cdot,\cdot\rangle$ to reflect this. As discussed in [5], see also [1], this normalization of the vertex is the most natural for 3d mirror symmetry.
@@
While this paper was in the final stages of preparation, we learned about a work in-progress by H. Nakajima, A. Okounkov, and Z. Zhou which contains a proof of Conjecture 1.3 for ADE quiver varieties (with no assumption about minuscule framings) using the universal deformation of quiver varieties [43]. The slant sum of two Dynkin quivers is almost never a Dynkin quiver; so the main interest of the techniques developed here lie outside ADE type.

### v2

@@ Abstract
We define the slant sum of quiver gauge theories, a gluing on the underlying quivers that identifies a gauge vertex with a framing vertex. Under some mild assumptions, we relate torus fixed points on the corresponding Higgs branches, which are Nakajima quiver varieties. Then we prove a “branching rule” relating the quasimap vertex functions before and after a slant sum and deduce a number of “factorization” corollaries.
Our construction is motivated by a factorization conjecture for the vertex functions of zero-dimensional quiver varieties, which can be approached inductively using the branching rule. In special cases, it also shows that vertex functions can be written as sums over reverse plane partitions, even outside ADE type.
We make some conjectures for Coulomb branches reflecting what can be seen on the Higgs side and prove them in ADE type. In particular, we obtain refined character formulas for the so-called “extremal” irreducible modules over shifted Yangians. We also study slant sums of Coulomb branches and their quantizations. We observe that for one-dimensional framing, the slant sum of Coulomb branches is the same as the product.
@@ 1.1 Slant sums of heaps
The slant sum of two posets was defined by Proctor in [undefav] to break certain posets into simpler, slant irreducible pieces. Some particular posets to which this applies are those that appear as heaps $H(w)$ of fully commutative elements $w$ of a Weyl group $W$, see [undefaw]. Such posets are defined in a straightforward combinatorial way from a choice of reduced word for $w$.
It was noticed in [undeft] that the heap of a minuscule element gives rise to a module over the preprojective algebra of the corresponding Dynkin quiver. Equivalently, it gives rise to a point in a Nakajima quiver variety. When $w$ is dominant minuscule, the corresponding quiver variety is a single point, and the preprojective algebra module studied in [undeft] is a choice of representative of this point.
Suppose $w$, $w^{(1)}$, and $w^{(2)}$ are dominant minuscule elements of some Weyl groups and let $\mathcal{M}$, $\mathcal{M}^{(1)}$, and $\mathcal{M}^{(2)}$ be the respective zero-dimensional Nakajima quiver varieties. If the heap of $w$ decomposes as a slant sum of the other two heaps, written $H(w)=H(w^{(1)})\#H(w^{(2)})$, then we call $\mathcal{M}$ the slant sum of the two other quiver varieties.
The aim of this paper is to revisit slant sums from the perspective of quiver varieties, or more precisely, quiver gauge theories. If the numerical data used to construct $\mathcal{M}^{(1)}$ and $\mathcal{M}^{(2)}$, which now need not be zero-dimensional, is compatible, we define a new quiver variety $\mathcal{M}:=\mathcal{M}^{(1)}\#\mathcal{M}^{(2)}$ which we call the slant sum of the two constituents. Unlike in [undeft], our procedure quickly leaves the world of Dynkin quivers, which we view as a feature.
@@ 1.2 Slant sums of quiver gauge theories
Let $Q^{(r)}$, $\mathsf{v}^{(r)}$, and $\mathsf{w}^{(r)}$ for $r\in\{1,2\}$ be two such collections. Suppose there exist vertices $\star_{1}\in Q^{(1)}_{0}$ and $\star_{2}\in Q^{(2)}_{0}$ such that $\mathsf{v}^{(1)}_{\star_{1}}=\mathsf{w}^{(2)}_{\star_{2}}$. We say that the two quiver gauge theories are slant summable, and we define their slant sum to be the quiver gauge theory defined by identifying the gauge vertex at $\star_{1}$ with the framing vertex at $\star_{2}$. We will use the symbol $_{\star_{1}}\#_{\star_{2}}$, or just #\#, throughout to denote the effect of this operation on various objects. For example, the quiver for the slant sum of the two quiver gauge theories is $Q=Q^{(1)}{}_{\star_{1}}\#_{\star_{2}}Q^{(2)}$.
@@
In the setting of the previous proposition, we will also denote $p^{(1)}\#p^{(2)}:=\Psi_{p^{(1)}}(p^{(2)})$. Theorem 1.1 generalizes the classical slant sum of heaps studied in [undefav], see also [undeft].
@@ 1.3 Slant sums and vertex functions
Our constructions also relate to certain quasimap counts. In the $K$-theoretic enumerative geometry of quiver varieties, the descendant vertex function is a key object [undefau]. It appears in Okounkov’s enumerative 3d mirror symmetry [undef, undefat], see also [undefe, undefn, undefo]. It is defined as the generating function of equivariant counts of quasimaps from $\mathbb{P}^{1}$ to $\mathcal{M}$ with a nonsingular condition at $\infty\in\mathbb{P}^{1}$. We denote it by
@@
and it is natural to wonder if there is any relationship between $V^{(\tau)}_{p}$, $V^{(\tau_{1})}_{p^{(1)}}$, and $V^{(\tau_{2})}_{p^{(2)}}$ when $\tau=\iota^{*}(\tau_{1}\otimes\tau_{2})$. This is the content of our next theorem.
@@
Under the hypotheses of Theorem 1.1, let $\tau=\iota^{*}(\tau_{1}\otimes\tau_{2})\in K_{\mathsf{T}\times G_{\mathsf{v}}}(\mathrm{pt})$. Then
@@
We obtain several corollaries in Section 3.4 in which Theorem 3.2 can be made more explicit.
Corollary 1.3 (Corollary 3.5).
If $\mathcal{M}^{(2)}=p^{(2)}$ and $\tau_{2}=1$, or more generally under the assumptions of Corollary 3.5, we have
where $z_{1}^{\prime}$ stands for a shift of $z_{1}$ by some powers of $z_{2}$.
Let $\hbar$ be the weight of the symplectic form on $\mathcal{M}$ and let $q$ be the equivariant parameters for the action of $\mathbb{C}^{\times}_{q}$ on $\mathbb{P}^{1}$. Under the Calabi-Yau specialization $\hbar=q$, Theorem 1.2 significantly simplifies.
Corollary 1.4 (Corollary 3.7).
Under the assumptions of Corollary 3.7
Theorem 1.2 also subsumes an additional formula present in the literature, the branching rule for the “nonstationary Ruijsenaars function”, studied, for example, in [undefk, undefaj, undefas, undefax]. We explain the details in Section 4.2.
1.4 Factorization for limits of vertex functions
Theorem 1.2 was motivated by 3d mirror symmetry, a proposed duality between Higgs and Coulomb branches [undefw, undefz, undefac, undefaaa]. Let $\mathsf{A}\subset\mathsf{T}$ be the subtorus preserving the symplectic form and let $\nu:\mathbb{C}^{\times}\to\mathsf{A}$. The cocharacter $\nu$ defines a partial resolution $\mathcal{M}^{!}$ of the corresponding Coulomb branch [undefj]. The GIT stability parameter $\theta$ used in the definition of $\mathcal{M}$ gives a cocharacter of a torus acting on $\mathcal{M}^{!}$. 3d mirror symmetry predicts that isolated points in $\mathcal{M}^{\nu(\mathbb{C}^{\times})}$ are in bijection with nonsingular points in $(\mathcal{M}^{!})^{\theta(\mathbb{C}^{\times})}$.
A conjecture of Smirnov and the first author [undefr] claims that $\lim_{a\to 0}\nu^{*}(V_{p})$, where $a$ is the coordinate on $\mathbb{C}^{\times}$, is a product of $q$-binomial series, one for each repelling weight of $T_{p^{!}}\mathcal{M}^{!}$. This implies a similar statement for the more extreme Calabi-Yau specialization $V_{p}|_{q=\hbar}$, in which $q$-binomials are replaced by ordinary binomials. Both of these are degenerations of a stronger statement, proven for type $A$ bow varieties in [undefe], relating vertex functions of $\mathcal{M}$ with those of $\mathcal{M}^{!}$. Conveniently, the degenerations can be checked even when there is no proposed definition of the vertex function of $\mathcal{M}^{!}$, which is currently the case outside in type $A$.
This conjecture reduces to the case where $\mathcal{M}$ is a single point as follows. An isolated fixed point $p\in\mathcal{M}^{\nu(\mathbb{C}^{\times})}$ can be viewed as the Higgs branch of its own quiver gauge theory, in which case $\lim_{a\to 0}\nu^{*}(V_{p})$ is equal to the vertex function of this zero-dimensional Higgs branch. Assuming a conjecture of Hilburn [undefae], Corollary 5.4 explains the similar reduction for Coulomb branches, which we verify for finite-type Dynkin quivers at the level of tangent space characters in (30) (generalizing [undefaf, Lemma 4.4] and [undefai, Equation (5.8)]).
1.5 Factorization for 0-dimensional vertex functions
@@
Now fix a quiver $Q$ without edge loops and a zero-dimensional quiver variety $\mathcal{M}:=\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})$ where the stabilty parameter is $\theta=(1,1,\ldots,1)$. We define coweights
$\lambda:=\sum_{i\in Q_{0}}\mathsf{w}_{i}\varpi_{i}^{\vee},\quad\mu:=\lambda-\sum_{i\in Q_{0}}\mathsf{v}_{i}\alpha_{i}^{\vee}$ (2)
@@
of the Kac-Moody Lie algebra $\mathfrak{g}_{Q}$ associated to the quiver $Q$. Here, $\varpi_{i}^{\vee}$ and $\alpha_{i}^{\vee}$ denote the fundamental coweights and simple coroots, respectively. Let $\Phi^{\pm,\text{re}}$ be the set of real positive or negative roots. Let $\Phi^{-,\text{re}}_{\mu}=\{\alpha\in\Phi^{-,\text{re}}\,\mid\,\langle\alpha,\mu\rangle>0\}$. It is a standard fact, see Lemma 7.4, that $\dim\mathcal{M}=0\iff\mu\in W\lambda$.
@@
The vertex function depends on a choice of polarization, and we choose the canonical polarization [^2] That is determined by the choice of orientation of the edges in the underlying graph of the quiver. Equivalently, we choose the polarization induced by the zero section of the cotangent bundle appearing in the symplectic reduction defining $\mathcal{M}$.. Then we define
@@
$\mathsf{V}_{\mathcal{M}}(z)=V_{\mathcal{M}}(z)_{z_{i}=z_{i}(-\hbar^{-1/2})^{a_{i}}},\quad\text{where}\quad a_{i}=\sum_{\begin{subarray}{c}e\\ h(e)=i\end{subarray}}\mathsf{v}_{t(e)}-\sum_{\begin{subarray}{c}e\\ t(e)=i\end{subarray}}\mathsf{v}_{h(e)}+\mathsf{w}_{i}$ (3)
@@
Let $A_{Q}=(a_{i,j})_{i,j\in Q_{0}}$ be the adjacency matrix of Q, i.e., $a_{i,j}$ is the number of arrows from vertex $i$ to vertex $j$. We identify the Kähler parameters with formal exponentials of simple roots via $e^{\alpha_{i}}=(\hbar/q)^{(\mathsf{v}-A_{Q}\cdot\mathsf{v})_{i}}z_{i}$. More explicitly,
$e^{\alpha_{i}}=z_{i}(\hbar/q)^{b_{i}},\quad b_{i}=\mathsf{v}_{i}-\sum_{\begin{subarray}{c}e\\ t(e)=i\end{subarray}}\mathsf{v}_{h(e)}$
Let $\Phi(x):=\prod_{i=0}^{\infty}(1-xq^{i})$. As discussed in [undef] and [undefe], the normalization of the vertex $\mathsf{V}_{\mathcal{M}}(z^{-1})$ is the most natural for 3d mirror symmetry.
@@
In type $A$, this conjecture follows from [undefe], though more direct proofs are available in special cases, see [undefo] and [undefs]. It was also proven by Jang and the first author for type $D$ quiver varieties with a single minuscule framing [undefp].
@@
While this paper was in the final stages of preparation, we learned about a work in-progress by H. Nakajima, A. Okounkov, and Z. Zhou which contains a proof of Conjecture 1.5 for ADE quiver varieties [undefaab]. The slant sum of two Dynkin quivers is almost never a Dynkin quiver; so the main interest of the techniques developed here lie outside ADE types.